# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

### v1

@@ Appendix B Bounds of delay errors
$\|\tilde{\delta}^{(m)}_{k}\|=$ $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|,$ (18)
@@
Lemma 4 (Delay error of signals).
Under Assumption 24, the difference between the latest and stale signals are bounded, i.e., $\forall m\in[M+1]$
$\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|\leq$ $\ C_{x}^{(m)}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (19)
$\|{\tilde{z}}^{(m-1)}_{k}-z^{(m)}_{k}\|\leq$ $\ C_{z}^{(m)}\sum_{m^{\prime}=1}^{m}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (20)
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq$ $\ C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|,$ (21)
where the constants above are defined by
$C_{x}^{(m)}:=$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m-1},$ (22)
$C_{z}^{(m)}:=$ $\ (L_{g,0}W_{\max,2})^{m-1},$ (23)
Proof of Lemma 4.
We first bound the delay of forward signal ${\tilde{x}}^{(m)}_{k}$. At the first layer, there is no delay, i.e. ${\tilde{x}}^{(1)}_{k}=x^{(1)}_{k}=x_{k}$. For other layers $m>1$
$=$ $\ L_{g,0}\|W^{(m)}_{k-(M-m)}{\tilde{x}}^{(m)}_{k}-W^{(m)}_{k}x^{(m)}_{k}\|$
where (a) follows the triangle inequality and the consistence of $\ell_{2}$-matrix norm, i.e. $\|Wu\|\leq\|W\|_{2}\|u\|$ where $W$ and $u$ are arbitrary matrix and vector, respectively. Using triangle inequality over the first term of RHS in (25), we have
and consequently,
$\ \|{\tilde{x}}^{(m+1)}_{k}-x^{(m+1)}_{k}\|$ (27)
$\leq$ $\ L_{g,0}(L_{g,0}W_{\max,2})^{m}\sum_{k^{\prime}=k-(M-m)}^{k-1}\|\Delta^{(m)}_{k^{\prime}}\|+L_{g,0}W_{\max,2}\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
$\ \|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|$ (28)
where the last inequality reindexes $m^{\prime}$ in the first summation.
Noticing that (25)-(28) also implies that
$\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|$ (31)
$\leq$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
$\ +W_{\max,2}(L_{g,0}W_{\max,2})^{M-m-1}L_{\ell,0}L_{g,0}L_{g,1}(L_{g,0}W_{\max,2})^{m-1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$=$ $\ L_{g,0}W_{\max,2}\|\delta^{(m+1)}_{k}-\tilde{\delta}^{(m+1)}_{k}\|$
$\ +(L_{g,0}W_{\max,2})^{M-1}L_{\ell,0}L_{g,1}\sum_{m^{\prime}=1}^{m-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
$\ \|\delta^{(M)}_{k}-\tilde{\delta}^{(M)}_{k}\|$ (32)
$\leq$ $\ (L_{\ell,1}L_{g,0}^{2}+L_{\ell,0}L_{g,1})(L_{g,0}W_{\max,2})^{M-1}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$
where (a) comes from Assumption 3 and Assumption 4.
$\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{m^{\prime\prime}-1}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (33)
$\leq$ $\ \sum_{m^{\prime\prime}=m}^{M-1}(L_{g,0}W_{\max,2})^{M-1-m^{\prime\prime}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (34)
$=$ $\ \frac{1-(L_{g,0}W_{\max,2})^{M-m}}{1-L_{g,0}W_{\max,2}}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|.$ (35)
Substituting (32) and (33) back into (31) we have
$\|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\leq C_{\delta}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|$ (36)
where the constant $C_{\delta}^{(m)}$ is defined by
This completes the proof. ∎
Lemma 5 (delay error of gradient).
Suppose Assumption 24 hold. Denote the stale gradient by
~$\tilde{\nabla}_{W^{(m)}}^{k}f:=\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}.$ (38)
The gradient delay is bounded by
~$\|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|.$ (39)
where the constant $C_{f}^{(m)}$ is defined by
Consequently, the norm square of the gradient delay is bounded by
~$\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}=$ $\ C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (41)
where the constant $C_{f}$ is defined by
$C_{f}:=M\sqrt{\sum_{m=1}^{M}(C_{f}^{(m)})^{2}}.$ (42)
Proof of Lemma 5.
By definition, the delay error is bounded by
~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (43)
$=$ $\ \|\tilde{\delta}^{(m)}_{k}\otimes{\tilde{x}}^{(m)}_{k}-\delta^{(m)}_{k}\otimes x^{(m)}_{k}\|$
$=$ $\ \|(\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k})\otimes{\tilde{x}}^{(m)}_{k}+\delta^{(m)}_{k}\otimes({\tilde{x}}^{(m)}_{k}-x^{(m)}_{k})\|$
$\leq$ $\ \|\tilde{\delta}^{(m)}_{k}-\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}\|+\|\delta^{(m)}_{k}\|\|{\tilde{x}}^{(m)}_{k}-x^{(m)}_{k}\|.$
According to Lemma 3 and Lemma 4, it follows that
~$\ \|\tilde{\nabla}_{W^{(m)}}^{k}f-\nabla_{W^{(m)}}f(W_{k};\xi_{k})\|$ (44)
$\leq$ $\ ((L_{g,0}W_{\max,2})^{m}C_{\delta}^{(m)}+(L_{g,0}W_{\max,2})^{M-m}L_{\ell,0}L_{g,0}C_{x}^{(m)})\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ C_{f}^{(m)}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
It completes the proof. ∎
@@ Appendix C Smoothness of the object
This section proves Lemma 1, i.e. the object is $L$-smooth. Before that, we introduce a lemma as the step stone.
Since the definition of function smoothness involves two weights $W$, $W^{\prime}$ and both of them generate a sequence of forward and backward signals, we require symbols to represent them. Using slightly abusing notation in this section, we denote the forward signals computed by $W$ and ${\tilde{W}}$ as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
Lemma 6 (Signal stability).
@@
$\ \|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}$ (52)
$\leq$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}^{(m-1)}-W^{(m-1)}\|^{2}+2L_{g,0}^{2}W_{\max,2}^{2}\|{\tilde{x}}^{(m-1)}-x^{(m-1)}\|^{2}$
@@
$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (53)
@@
$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq C_{\delta^{\prime}}^{(m)}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$ (61)
@@
Analogous to the proof of Lemma (6), we use $x^{(m)}$/$\delta^{(m)}$ and ${\tilde{x}}^{(m)}$/$\tilde{\delta}^{(m)}$ to denote the forward/backward signals related to $W$ and ${\tilde{W}}$, respectively. It follows by definition of gradient (4) that
$\ \|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}$ (64)
$=$ $\ \sum_{m=1}^{M}\|\nabla_{W^{(m)}}f(W;\xi)-\nabla_{W^{(m)}}f({\tilde{W}};\xi)\|^{2}$ (65)
@@
Taking square root on both sides on (66) completes the proof. ∎
@@ Appendix D Proof of Theorem 1: Analog SGD with synchronous pipeline
¯$\bar{\nabla}f(W_{k};\xi_{k}):=\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b}),$ (68)
@@
$=$ $\ \frac{1}{B^{2}}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$
$\leq$ $\ \sigma^{2}/B.$
@@
where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
@@
$\ \frac{L_{f}}{2}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}\|^{2}]$ (72)
$\leq$ $\ L_{f}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$
@@
$\ {\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]$ (73)
$\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}+\frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}].$
According the dynamic of Analog SGD (5), the last term in the RHS of (73) is upper bounded by
@@
${\mathbb{E}}_{\xi_{k,b}}[\|f(W_{k};\xi_{k,b})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]$ (75)
$\leq$ $\ \|\nabla f(W_{k})\|^{2}+\sigma^{2}.$
@@
Organizing the terms, taking expectation, averaging (76) over $k$ from 00 to $K-1$, we obtain
@@
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max}^{2}/\tau^{2}}+\sigma^{2}S_{K}$
@@ Appendix E Proof of Theorem 2: Analog SGD with asynchronous pipeline
where the second inequality comes from the assumption that noise has expectation 00 (Assumption 1)
@@
By the dynamic of Analog SGD, the last in the RHS of (E) is bounded by
@@
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}M^{2}C_{f}^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
@@
${\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (86)
@@
$=$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
@@
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
@@
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
@@
${\mathbb{V}}_{k}:=$ $\ f(W_{k})-f^{*}$ (89)
$+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
Notice that the last term above has the following recursion from $k$ to $k+1$
$\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-(k+1)+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (90)
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k-1}(k^{\prime}-k+M+1)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$\ +(M-m+1)\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}.$
@@
$\ {\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]$ (91)
$=$ $\ f(W_{k+1})-f^{*}+\frac{\alpha C_{f}M^{2}}{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime}+1)}^{k}(k^{\prime}-{k+1}+M+1)\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime}-1)}_{k^{\prime}}\|^{2}]$
@@
$\ -\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime}+1)}^{k}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
By the definition of $\Delta^{(m)}_{k}$ and the dynamic of Analog SGD, we bound the last term of the RHS of (91) by
$\ \sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}=\sum_{m^{\prime}=1}^{M}\|W^{(m)}_{k+1}-W^{(m)}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}$ (92)

### v2

@@ Appendix B Useful Lemmas and their proof
$\|\tilde{\delta}^{(m)}_{k}\|$= $\ \|\tilde{\delta}^{(m+1)}_{k}W^{(m+1)}_{k}\nabla g^{(m)}({\tilde{z}}^{(m)}_{k})\|$ (59)
@@ Appendix C Smoothness of the object
This section proves Lemma 2, i.e. the object is $L$-smooth. Before that, we introduce a lemma as the step stone. Since the definition of function smoothness involves two weights $W$, $W^{\prime}$ and both of them generate a sequence of forward and backward signals, we require symbols to represent them. Using slightly abusing notation in this section, we denote the forward signals computed by $W$ and $\tilde{W}$} as $x^{(m)}$ and ${\tilde{x}}^{(m)}$, respectively, and denote the backward signals as $\delta^{(m)}$ and $\tilde{\delta}^{(m)}$, respectively.
Lemma 9 (Signal stability, restatement of Lemma 1).
@@
$\|{\tilde{x}}^{(m)}-x^{(m)}\|^{2}\leq$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\|{\tilde{W}}^{(m-1)}-W^{(m-1)}\|^{2}+2L_{g,0}^{2}W_{\max,2}^{2}\|{\tilde{x}}^{(m-1)}-x^{(m-1)}\|^{2}$ (67)
@@
$=$ $\ 2L_{g,0}^{2}(L_{g,0}W_{\max,2})^{2(m-1)}\sum_{m^{\prime}=1}^{m-1}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}$
@@
$\|\tilde{\delta}^{(m)}-\delta^{(m)}\|^{2}\leq C_{\delta^{\prime}}^{(m)}\sum_{m^{\prime}=1}^{M}\|{\tilde{W}}^{(m^{\prime})}-W^{(m^{\prime})}\|^{2}.$ (75)
@@
Analogous to the proof of Lemma (9), we use $x^{(m)}$$\delta^{(m)}$ and ${\tilde{x}}^{(m)}$$\tilde{\delta}^{(m)}$ to denote the forward/backward signals related to $W$ and $\tilde{W}$}, respectively. It follows by definition of gradient (5) that
$\|\nabla f(W;\xi)-\nabla f({\tilde{W}};\xi)\|^{2}=$ $\ \sum_{m=1}^{M}\|\nabla_{W^{(m)}}f(W;\xi)-\nabla_{W^{(m)}}f({\tilde{W}};\xi)\|^{2}$ (78)
@@
Taking the square root on both sides of (79) completes the proof. ∎
@@ Appendix D Proof of Theorem 1: Analog SGD with synchronous pipeline
$\bar{\nabla}f(W_{k};\xi_{k}):=\frac{1}{B}\sum_{b=1}^{B}\nabla f(W_{k};\xi_{k,b}),$ (81)
@@
$=$ $\ \frac{1}{B^{2}}\sum_{b=1}^{B}{\mathbb{E}}_{\xi_{k,b}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]\leq\sigma^{2}/B.$
@@
where the second inequality comes from the assumption that the expectation of the noise is 0 (Assumption 1)
@@
$\frac{L_{f}}{2}\mathbb{E}_{\varepsilon_{k}}[\|W_{k+1}-W_{k}\|^{2}]\leq$ $\ L_{f}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha(\bar{\nabla}f(W_{k};\xi_{k})-\nabla f(W_{k}))\|^{2}]$ (85)
@@
${\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[f(W_{k+1})]\leq$ $\ f(W_{k})-\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}+\frac{\alpha^{2}L_{f}\sigma^{2}}{B}$ (86)
$\ +\frac{1}{2\alpha}{\mathbb{E}}_{\{\xi_{k,b}:b\in[B]\}}[\|W_{k+1}-W_{k}+\alpha\bar{\nabla}f(W_{k};\xi_{k})\|^{2}].$ (87)
According the dynamics of Analog-SGD-SP, the last term in the RHS of (86) is bounded by
@@
${\mathbb{E}}_{\xi_{k,b}}[\|f(W_{k};\xi_{k,b})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k,b})-\nabla f(W_{k})\|^{2}]\leq\|\nabla f(W_{k})\|^{2}+\sigma^{2}.$ (89)
@@
Organizing the terms, taking expectation, averaging (90) over $k$ from 0 to K1K-1, we obtain
@@
$\leq$ $\ 4\sqrt{\frac{(f(W_{0})-f^{*})\sigma^{2}L_{f}}{BK}}\frac{1}{1-W_{\max,\infty}^{2}/\tau^{2}}+\sigma^{2}S_{K}$
@@ Appendix E Proof of Theorem 2: Analog SGD with asynchronous pipeline
where the second inequality comes from the assumption that the expectation of the noise is 0 (Assumption 1)
@@
By the dynamics of Analog SGD, the last in the RHS of (E) is bounded by
@@
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}\mathbb{E}_{\xi_{k}}[\|\tilde{\nabla}^{k}f-\nabla f(W_{k};\xi_{k})\|^{2}]$
$\leq$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}M^{2}C_{f}^{2}\mathbb{E}_{\xi_{k}}\left[\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}\right].$
@@
${\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})\|^{2}]=$ $\ \|\nabla f(W_{k})\|^{2}+{\mathbb{E}}_{\xi_{k}}[\|\nabla f(W_{k};\xi_{k})-\nabla f(W_{k})\|^{2}]$ (100)
@@
$=$ $\ \frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}]$
@@
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
@@
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\mathbb{E}_{\xi_{k}}\Big[\underbrace{\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}}_{=:\psi_{k}}\Big].$
To deal with $\psi_{k}$ in the RHS of (102), we define an auxiliary function
$\Psi_{k}:=\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (103)
which has the following recursion from $k$ to $k+1$
$\Psi_{k+1}=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-(k+1)+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$ (104)
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$=$ $\ \sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k-(M-m^{\prime})}^{k-1}(k^{\prime}-k+M)\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}-\sum_{m^{\prime}=1}^{M}\sum_{k^{\prime}=k+1-(M-m^{\prime})}^{k}\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}$
$\ +M\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}-\sum_{m^{\prime}=1}^{M}m^{\prime}\|\Delta^{(m^{\prime})}_{k-(M-m^{\prime})}\|^{2}$
$\leq$ $\ \Psi_{k}-\psi_{k}+M\|W_{k+1}-W_{k}\|^{2}$
where the last inequality holds by definition of $\Delta^{(m^{\prime})}_{k}$, i.e.,
$\sum_{m^{\prime}=1}^{M}\|\Delta^{(m^{\prime})}_{k}\|^{2}=\sum_{m^{\prime}=1}^{M}\|W^{(m^{\prime})}_{k+1}-W^{(m^{\prime})}_{k}\|^{2}=\|W_{k+1}-W_{k}\|^{2}.$ (105)
@@
${\mathbb{V}}_{k}:=f(W_{k})-f^{*}+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\Psi_{k}.$ (106)
@@
${\mathbb{E}}_{\xi_{k}}[{\mathbb{V}}_{k+1}]=$ $\ f(W_{k+1})-f^{*}+\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{2}\Psi_{k+1}$ (107)
@@
$\ +\frac{\alpha}{2}(1+\frac{1}{u})(1+W_{\max,\infty}/\tau)^{2}C_{f}^{2}M^{3}\sum_{m^{\prime}=1}^{M}\mathbb{E}_{\xi_{k}}[\|\Delta^{(m^{\prime})}_{k^{\prime}}\|^{2}].$
we bound the last term of the RHS of (107) by
$\ \|W_{k+1}-W_{k}\|^{2}$ (108)